OpenStudy (anonymous):

how would you graph (x-1)^2 + (y-1)^2+ (z-1)^2 < 1/4

myininaya (myininaya):

(1,1,1) is center of sphere

myininaya (myininaya):

if we wanted to graph (x-1)^2+(y-1)^2+(z-1)^2=1/4 then (1,1,1) is center and r=1/2 so draw a sphere with center (1,1,1) and radius 1/2 then shade everything inside the sphere since you want all spheres with center (1,1,1) and radi less than 1/2 and make the graph for (x-1)^2+(y-1)^2+(z-1)^2=1/4 broken (not a solid curve)

OpenStudy (anonymous):

i know it is an open set, but i am confused on how to draw this on a 3d graph lol

OpenStudy (anonymous):